# thanks: Contributed Equally

### v1

@@ Abstract
We develop an extended Dynamical Mean Field Theory framework to analyze gene regulatory networks (GRNs) incorporating epigenetic modifications. Building on the Hopfield network model analogy to spin glass systems, our approach introduces dynamic terms representing DNA methylation and histone modification to capture their regulatory influence on gene expression. The resulting formulation reduces high-dimensional GRN dynamics to effective stochastic equations, enabling the characterization of both stable and oscillatory states in epigenetically regulated systems. This framework provides a tractable and quantitative method for linking gene regulatory dynamics with epigenetic control, offering new theoretical insights into developmental processes and cell fate decisions.
@@ I Introduction
A well-defined mathematical modelling framework is essential for mapping real-world biological processes into mathematical representations, enabling the study of both local and global system behaviours [1]. Such an approach provides a systematic foundation for understanding complex biological mechanisms that cannot be fully explained through qualitative reasoning or experimental observation alone [2, 1].
In the context of computational biology–particularly developmental biology–experimental evidence has traditionally dominated over theoretical or simulation-based analyses [3, 4, 5, 6, 7]. However, to obtain robust and generalizable conclusions from experimental results, a mathematical or modelling framework is crucial [3, 4, 5]. This becomes especially important when the available experimental data are limited or when experiments are costly and labour-intensive. Biological experiments, such as those involving single-cell RNA sequencing (scRNA-seq), are excellent examples: they require substantial resources, careful experimental design, and stringent protocols [6, 7, 8, 2]. In such cases, mathematical models play a vital role in interpolating incomplete data, identifying consistent patterns, and extending biological insights beyond direct experimental reach.
Since its introduction, scRNA-seq has emerged as a powerful technique for studying developmental processes in both mouse and human systems [8, 9]. It enables researchers to investigate how cells transition through developmental stages by monitoring the continuous changes in gene expression profiles [6, 7]. However, experimentally capturing these continuous, time-dependent processes is extremely challenging and expensive. Gene regulatory network (GRN) models provide a valuable framework for describing and analysing the underlying dynamics of gene expression and for identifying the regulatory mechanisms that control cell fate decisions [10, 11, 12, 13, 4, 5, 2, 14].
Among various mathematical models used to describe GRNs, the Hopfield network model has gained considerable attention [15, 16, 17]. In this model, gene states are represented as binary variables—either “on” or “off”—depending on a defined threshold, drawing a close analogy to the classical Ising model from statistical physics [17]. This correspondence allows researchers to apply well-established analytical tools from spin glass theory to gene regulatory systems, subject to suitable approximations [17]. While both models admit similar energy functions, the Ising model is an equilibrium system defined by a Hamiltonian, whereas the Hopfield network uses an energy function as a Lyapunov function for dissipative dynamics, leading to different interpretations of steady states and dynamics [15, 16].
A complementary conceptual approach to studying gene regulatory dynamics is Waddington’s epigenetic landscape [14, 3, 5]. In this metaphor, cellular differentiation is visualized as a ball rolling down a landscape of valleys and ridges, where each valley represents a stable cell fate. Mathematically, this corresponds to constructing a potential landscape that captures the global dynamics of the system. However, for high-dimensional and non-linear GRNs, constructing such a potential analytically remains an open and challenging problem [6, 7, 14]. Only limited methods exist to approximate or infer these landscapes from experimental or simulated data [6, 7].
A promising theoretical framework for addressing these challenges is Dynamical Mean Field Theory (DMFT), originally developed by Sompolinsky and colleagues to study the collective dynamics of spin glass systems [18, 19, 20, 21, 22]. DMFT simplifies the analysis of high-dimensional, complex networks by transforming global interactions into effective local stochastic equations [18, 23]. This reduction provides a tractable means to study the macroscopic behaviour of large, non-linear systems, including dissipative biological networks [24, 23]. Applying DMFT to GRNs could thus offer valuable insights into their collective dynamics and the emergent properties of developmental systems, especially when the construction of explicit potential functions is analytically intractable [18].
The problem becomes even more intricate when epigenetic modifications are incorporated into GRN models [11, 12, 13, 10]. Epigenetic mechanisms, such as DNA methylation and histone modification, can dynamically reprogram gene expression without altering the underlying DNA sequence. These processes introduce additional layers of regulation that can lead to both stable and oscillatory cellular states, increasing the system’s dynamical complexity [11, 12, 13, 10]. Despite their biological significance, there is currently no well-established analytical framework to treat epigenetic modifications within the context of GRN dynamics. While some approximations can capture these effects qualitatively [6, 7], a comprehensive theoretical formulation remains lacking.
In this work, we aim to extend Sompolinsky’s idea to include epigenetic modification terms in GRN models. Our objective is to develop a comprehensive and reliable mathematical framework that can systematically analyse the complex dynamics of epigenetically regulated GRNs. This approach not only bridges the gap between experimental observations and theoretical analysis but also provides a ready-to-use tool for exploring developmental processes, cell fate decisions, and the emergence of diverse cellular states in biological systems.
@@ II Mechanistic model of GRN–epigenetic interactions
We turn to DMFT because it provides a powerful method—introduced by Sompolinsky–for reducing high-dimensional network dynamics to an effective one-dimensional description governed by Newtonian-like equations, thereby enabling tractable analysis of complex systems through local dynamics [18, 20, 21, 22, 19]. In our model, however, the presence of the epigenetic variable $\theta_{i}$ introduces an intrinsic feedback loop and slow timescale, complicating the direct use of conventional DMFT [18, 20, 21, 22, 19].
Thus, our goal is to extend and adapt the DMFT methodology to handle this feedback-driven structure. By developing a mean-field formulation that separates the dynamics of the interacting components and leverages timescale separation, we aim to partially decouple the two coupled ODEs and obtain a tractable theoretical understanding of the full homeorhesis process [11, 12, 13, 10].
@@ II.1 Mean-Field framework
This raises the natural question: how can DMFT be applied to this model, and how does the transformed local dynamics—and hence the global dynamics—reshape the potential landscape? Constructing such a landscape would provide insight into cellular differentiation under slow epigenetic regulation, which gradually reprograms gene-expression dynamics during development [11, 12, 13, 10].
@@ II.2 Disparity of characteristic timescales
Under the assumption $\nu\ll 1$, we observe a clear separation of timescales for $x_{i}$ and $\theta_{i}$. The remaining parameters are unaffected in this limit. As a consequence, we have $\dot{\theta}_{i}\approx 0$ when deriving the equation for $\dot{y}_{i}$. However, this does not imply that $\theta_{i}$ becomes time-independent. Rather, the final solution for $y_{i}$ still depends on the time-dependent $\theta_{i}$, but its time derivative is sufficiently small that the integration proceeds as if $\theta_{i}$ were approximately constant. Under this assumption, the governing equation for $y_{i}$ of the $i^{\text{th}}$ agent becomes
@@ III.2 Evaluating the expected noise contribution
In the presence of non-negligible noise, we must evaluate the contribution arising from the noise term. To do so, we generalize $X_{i}$ so that it is averaged not only over time $t$ but also over the noise realization $\eta$, which itself is time-dependent. In this context, we make use of the complex integral relation
@@
$\tilde{C}(t)=\int_{0}^{t}\exp(\nu-2t)\left[\int_{0}^{\nu}C_{\eta}(u)\mathrm{d}u\right]\mathrm{d}\nu.$
@@ V Autocorrelation Function and Time-Scale Separation in the DMFT Framework
A natural question then arises: to what extent are the network dynamics modified by the presence of a slow epigenetic modification term in a gene regulatory network, given that the prototype model without such a term has been extensively studied, beginning with Sompolinsky’s original work [18]. In particular, can the autocorrelation function still capture the essential effects of epigenetic reprogramming when slow epigenetic variables modulate the regulatory interactions? Furthermore, how is the effective Newtonian dynamics altered through changes in the autocorrelation function induced by epigenetic modifications? Specifically, how does the corresponding effective potential deform in the presence of such slow epigenetic feedback, and to what extent does it reshape the dynamical landscape of the system?
@@
In the next step, we define the atomic autocorrelation function $\Delta(\tau)=\mathbf{E}_{t,J}\left[y_{k}^{1}(t)y_{k}^{1}(t+\tau)\right]$. By substituting Eq. (6) into this definition, we obtain the following differential equation,
@@
By recognizing that $y^{(1)}(t)$ is driven by Gaussian noise—and is therefore not specific to the $k$-th agent—we may drop the index $k$. Since these variables are noise-driven, they are themselves Gaussian random variables, fully characterized by their moments,
$\mathbf{E}_{t,J}\left[y^{1}(t)\right]=\mathbf{E}_{t,J}\left[y^{1}(t+\tau)\right]=0,$
$\mathbf{E}_{t,J}\left[y^{1}(t)y^{1}(t+\tau)\right]=\Delta(\tau).$
@@
$y^{1}(t)$ $=\hat{\alpha}(\tau)z_{1}+\hat{\beta}(\tau)z_{3},$
$y^{1}(t+\tau)$ $=\hat{\alpha}(\tau)z_{2}+\hat{\gamma}(\tau)z_{3},$
@@
where we used $y(t)=y^{0}(t)+y^{1}(t)$ and $\mathrm{D}z_{j}=\frac{1}{\sqrt{2\pi}}\exp(-z_{j}^{2}/2)\mathrm{d}z_{j}$.
@@
By slightly redefining the parameters—factoring $g$ out of the variance and incorporating it into the gain parameter $\beta$—we obtain the dimensionless effective gain $\beta g$. This rescaling does not alter the system dynamics but simplifies the subsequent derivations. Accordingly, we can define an effective potential $V^{t}(\Delta)$ of the form:
@@ V.1 Viable Autocorrelation States
The system’s autocorrelation function $\Delta$ is subject to several constraints: boundedness ($\Delta_{0}\geq|\Delta|$), evenness ($\Delta(\tau)=\Delta(-\tau)$), and no initial momentum ($\dot{\Delta}(\tau=0)=0$). Consequently, the following regimes are discussed only for $\Delta_{0}\leq\Delta\leq\Delta_{0}$, which defines the most rudimentary valid range of $\Delta$. For $\theta=0=c$, we recover the same potential and derivations as in Sompolinsky’s work. In this section, we focus solely on the viability of certain autocorrelation functions. The stability of these states—while related to the potential—cannot be directly inferred from its quantitative behaviour and will be discussed separately.
@@
The second regime arises for $\beta g>1$ and $\Delta_{0}<\tilde{\Delta}_{1}$ for some specific value $\tilde{\Delta}_{1}$, as illustrated in Fig. 3. For a given $V(\Delta)$ with $\theta=0$, there exists a continuum of self-consistent solutions of Eq. (19), corresponding to periodic orbits and limited cycles. Non-zero values of $\theta$ distort and shift the potential, introducing asymmetry. This asymmetry eliminates some previously allowed limiting cycles, since they would violate the boundedness constraint ($\Delta>\Delta_{0}$). The only remaining allowed solution is the potential minimum, which represents a non-zero fixed point of the system—a spin glass freezing state. The location of this fixed point can be tuned by adjusting $\theta$, offering a way to control the system’s behavior.
@@ VII Summary and Discussion
In our approach, we have merged the classical problem of constructing a potential well with the Waddington metaphor, extending it to a slightly deformed version when various perturbations—such as feedback mechanisms—are present in the network [10, 12, 11, 13]. These perturbations are key to analysing the stability of fixed points and the nature of bifurcations, which are critical for understanding cellular differentiation processes and disease progression, such as cancer metastasis [31, 32]. Our study begins with this motivation, focusing on the construction of a potential landscape when an epigenetic feedback term is included in the GRN. This feedback term is typically used to describe the reprogramming process in GRNs, which is crucial for cellular reprogramming or repair [10, 12, 11, 13].
@@
The complexity arises when perturbations or feedback terms are introduced into the system, adding extra dimensions to the biophysical model. This makes the construction of the potential landscape more challenging, as the additional dimensions complicate the mathematical analysis. In this work, our main motivation was to provide a readily available mathematical mechanism to handle these feedback terms, which describe reprogramming in GRNs, to construct the potential surface or landscape. This approach allows us to analyse the stability of global dynamics in such systems.
We also aimed to compare the deformation of the potential landscape caused by the feedback term with the non-feedback case through proper mathematical analysis using the traditional DMFT framework. The goal was to provide a mechanism that would save time, bypassing the need to generate hypotheses from large datasets by directly predicting a mathematical model from data using modern machine learning algorithms. The primary focus of our DMFT analysis was to observe how the autocorrelation function and the matching moments were affected by varying the strength of the feedback term. Our results indicate that even small changes in the feedback term can lead to significant deformations in the potential landscape.
While our method has certain limitations—given the mathematical complexity and the length of the analysis—it provides a valuable framework for constructing a reliable version of the potential landscape of GRNs with feedback. This approach allows us to analyse the strength of feedback and its effect on the potential surface, providing insight into various intrinsic parameters in both prototype and real models.

### v2

@@ Abstract
Understanding how gene regulatory networks (GRNs) give rise to stable and dynamic cellular states remains a central challenge in theoretical biology, particularly when slow epigenetic feedback reshapes the underlying regulatory landscape. While experimental approaches such as single-cell transcriptomics reveal rich dynamical behaviour, a tractable theoretical framework that links gene expression, epigenetic control, and collective dynamics remains challenging. Here, we develop an extended Dynamical Mean Field Theory (DMFT) framework for GRNs that incorporates epigenetic modifications as slow, feedback-driven variables. Building on the analogy between Hopfield networks and spin glass systems, we derive effective stochastic equations that reduce high-dimensional dynamics to a tractable form across multiple timescales. This formulation enables quantitative characterization of both stable and oscillatory regimes and reveals how epigenetic feedback reshapes the effective potential landscape governing cell fate decisions. Our model shows how epigenetic feedback regulation dynamically reshapes the Waddington landscape. Our results and methodology provide a unified theoretical framework for understanding developmental dynamics and epigenetic reprogramming in complex biological systems.
@@ I Introduction
Understanding how high-dimensional biological systems generate stable yet flexible dynamical behavior is a central challenge at the interface of physics and biology [1, 2, 3, 4]. Gene regulatory networks (GRNs) govern cellular identity by orchestrating interactions among transcription factors, chromatin states, and signaling pathways [5, 1, 6, 7]. Yet, their dynamics are shaped not only by fast transcriptional processes but also by slower, history-dependent epigenetic modifications [6, 8]. While experimental advances such as single-cell and spatial transcriptomics have revealed rich dynamical heterogeneity across developmental and disease contexts [9, 10, 11, 12, 13, 14, 15], a general theoretical framework that links gene expression dynamics to epigenetic regulation across timescales remains lacking [8, 16].
Classical approaches to GRN dynamics have drawn heavily on analogies to statistical physics, particularly through the connection between recurrent neural networks and spin-glass systems [17, 2, 13]. In these models, stable gene expression patterns correspond to attractors of an energy-like landscape, providing an appealing interpretation of cell types as minima of an effective potential [1, 13]. However, such formulations typically assume symmetric interactions and equilibrium dynamics, thereby restricting the system to gradient flows that converge to fixed points [3, 18, 4]. This limitation is at odds with a growing body of evidence that biological systems operate far from equilibrium, exhibiting oscillations, lineage transitions, and metastable states that cannot be captured within purely dissipative frameworks[5, 10, 19, 16, 9, 14, 15].
At the same time, epigenetic regulation introduces an additional layer of complexity by modulating the effective regulatory interactions over longer timescales [7, 6, 8]. Chromatin accessibility, histone modifications, and DNA methylation dynamically reshape the regulatory landscape, enabling memory, hysteresis, and context-dependent responses [6, 8, 16, 20]. These processes introduce feedback between gene expression and the underlying regulatory architecture, effectively coupling fast transcriptional dynamics to slow structural changes [6, 21, 8, 16]. Despite their central role in development and reprogramming, such feedback mechanisms remain difficult to incorporate into existing theoretical models, which often treat regulatory interactions as static or assume a separation between dynamics and structure.
To address this gap, we develop a Dynamical Mean Field Theory (DMFT) [22, 23, 24] that encompasses GRNs with epigenetic feedback. DMFT provides a powerful approach for analyzing high-dimensional interacting systems by reducing their collective behavior to an effective stochastic process that captures both fluctuations and self-consistent interactions [22, 23, 25, 24, 26]. Originally developed in the context of disordered systems and spin glasses [22, 23], and later applied to neural network dynamics [27, 25], DMFT is particularly well suited to studying large, heterogeneous networks with random or partially structured connectivity. Here, we generalize DMFT to incorporate slow, feedback-driven modifications of the interaction matrix, thereby coupling fast gene-expression dynamics to evolving epigenetic states.
Within this formulation, each gene is described by an effective stochastic differential equation whose statistics are determined self-consistently by the network [22, 23, 25, 27]. Epigenetic variables evolve on a slower timescale and modulate the strength and structure of regulatory interactions [8, 16], thereby reshaping the effective potential landscape. This leads to a coupled system in which the landscape itself is no longer fixed but adapts to the system’s trajectory [8, 16]. As a result, the dynamics cannot be described solely in terms of convergence to static attractors; instead, they exhibit a rich interplay among stability, fluctuations, and the slow drift of the underlying regulatory structure.
Our analysis reveals how epigenetic feedback qualitatively alters the dynamical regimes of GRNs. In particular, we identify conditions under which stable attractors persist and regimes in which feedback induces oscillatory or metastable behavior. The framework allows us to characterize how memory and hysteresis emerge from the coupling between fast and slow variables, and how transitions between cellular states can be driven by gradual reshaping of the regulatory landscape rather than by external perturbations alone. Importantly, the DMFT reduction provides tractable equations for these phenomena, enabling analytical insight into systems that would otherwise be intractable due to their high dimensionality.
More broadly, this work establishes a unified theoretical perspective on gene regulation as a non-equilibrium dynamical system with evolving interactions. By integrating epigenetic feedback into a mean-field description, we bridge the gap between energy landscape models and time-dependent regulatory architectures, providing a framework that connects statistical physics, dynamical systems theory, and modern high-throughput biological data. This approach opens new avenues for understanding developmental processes, cellular plasticity, and epigenetic reprogramming in complex biological systems.
@@ II Mechanistic model of GRN–epigenetic interactions
We turn to DMFT because it provides a powerful method–introduced by Sompolinsky–for reducing high-dimensional network dynamics to an effective one-dimensional description governed by Newtonian-like equations, thereby enabling tractable analysis of complex systems through local dynamics [22, 24, 25, 27, 23]. In our model, however, the presence of the epigenetic variable $\theta_{i}$ introduces an intrinsic feedback loop and slow timescale, complicating the direct use of conventional DMFT [22, 24, 25, 27, 23].
Thus, we extend and adapt the DMFT methodology to handle this feedback-driven structure. By developing a mean-field formulation that separates the dynamics of the interacting components and leverages timescale separation, we aim to partially decouple the two coupled ODEs and obtain a tractable theoretical understanding of the full homeorhesis process [16, 6, 21, 8].
@@ II.1 Mean-Field framework
This raises the natural question: how can DMFT be applied to this model, and how does the transformed local dynamics–and hence the global dynamics–reshape the potential landscape? Constructing such a landscape would provide insight into cellular differentiation under slow epigenetic regulation, which gradually reprograms gene-expression dynamics during development [16, 6, 21, 8].
@@ II.2 Disparity of characteristic timescales
Under the assumption $\nu\ll 1$, we observe a clear separation of timescales for $x_{i}$ and $\theta_{i}$. The remaining parameters are unaffected in this limit. As a consequence, we have $\dot{\theta}_{i}\approx 0$ when deriving the equation for $\dot{y}_{i}$. Importantly, this does not mean that $\theta_{i}$ is truly time-independent: the final solution for $y_{i}$ still depends on the full, slowly varying $\theta_{i}$. Depending on the timescale of interest, one can therefore either treat $\theta_{i}$ as effectively constant (fast dynamics of $x_{i}$) or account for its slow temporal evolution (longer timescales). Which approximation is appropriate should be clear from the context, based on the current time-scale of interest. Under this assumption, the governing equation for $y_{i}$ of the $i^{\text{th}}$ agent becomes
@@ III.2 Evaluating the expected noise contribution
In the presence of non-negligible noise, we must evaluate the contribution arising from the noise term. To do so, we generalize $X_{i}$ so that it is averaged not only over time $t$ but also over the noise realization $\eta$, which itself is time-dependent. This can be interpreted as treating the coupling of a given agent to the rest of the system, and hence as accounting for the influence of the agent’s surroundings. In this context, we make use of the complex integral relation
@@
Notice that for Gaussian noise with zero mean and bounded, possibly time-dependent variance, the noise average produces a bounded analytic contribution. This allows us to perform the noise and time averages on the regularized integrand at fixed $\epsilon>0$. The exchange of the noise and time averages with the $u$-integral is physically justified by the damping provided by the noise average.
@@
$\tilde{C}(t)=\int_{0}^{t}\exp(\tau-2t)\left[\int_{0}^{\tau}C_{\eta}(s)\mathrm{d}s\right]\mathrm{d}\tau.$
@@ V Autocorrelation Function and Time-Scale Separation in the DMFT Framework
A natural question then arises: to what extent are the network dynamics modified by the presence of a slow epigenetic modification term in a GRN, given that the prototype model without such a term has been extensively studied, beginning with Sompolinsky’s original work [22]. In particular, can the autocorrelation function still capture the essential effects of epigenetic reprogramming when slow epigenetic variables modulate the regulatory interactions? Furthermore, how is the effective Newtonian dynamics altered through changes in the autocorrelation function induced by epigenetic modifications? Specifically, how does the corresponding effective potential deform in the presence of such slow epigenetic feedback, and to what extent does it reshape the dynamical landscape of the system?
@@
In the next step, we define the atomic autocorrelation function $\Delta(\tau)=\mathbf{E}_{t,J}\left[y_{k,1}(t)y_{k,1}(t+\tau)\right]$. By substituting Eq. (6) into this definition, we obtain the following differential equation,
@@
By recognizing that $y_{k,1}(t)$ is driven by Gaussian noise—and is therefore not specific to the $k$-th agent—we may drop the index $k$. Since these variables are noise-driven, they are themselves Gaussian random variables, fully characterized by their moments,
$\mathbf{E}_{t,J}\left[y_{1}(t)\right]=\mathbf{E}_{t,J}\left[y_{1}(t+\tau)\right]=0,$
$\mathbf{E}_{t,J}\left[y_{1}(t)y_{1}(t+\tau)\right]=\Delta(\tau).$
@@
$y_{1}(t)$ $=\hat{\alpha}(\tau)z_{1}+\hat{\beta}(\tau)z_{3},$
$y_{1}(t+\tau)$ $=\hat{\alpha}(\tau)z_{2}+\hat{\gamma}(\tau)z_{3},$
@@
where we used $y_{i}(t)=y_{i,0}(t)+y_{i,1}(t)\equiv y_{i,0}(t)+y_{1}(t)$ and $\mathrm{D}z_{j}=\frac{1}{\sqrt{2\pi}}\exp(-z_{j}^{2}/2)\mathrm{d}z_{j}$.
@@
By slightly redefining the parameters–factoring $g$ out of the variance and incorporating it into the gain parameter $\beta$–we obtain the dimensionless effective gain $\beta g$. This rescaling does not alter the system dynamics but simplifies the subsequent derivations. Accordingly, we can define an effective potential $V^{t}(\Delta)$ of the form:
@@ V.1 Viable Autocorrelation States
The system’s autocorrelation function $\Delta$ is subject to several constraints: boundedness ($\Delta_{0}\geq|\Delta|$), evenness ($\Delta(\tau)=\Delta(-\tau)$), and no initial momentum ($\dot{\Delta}(\tau=0)=0$). Consequently, the following regimes are discussed only for $\Delta_{0}\leq\Delta\leq\Delta_{0}$, which defines the most rudimentary valid range of $\Delta$. For $\theta=0=c$, we recover the same potential and derivations as in Sompolinsky’s work. In this section, we focus solely on the viability of certain autocorrelation functions. The stability of these states–while related to the potential–cannot be directly inferred from its quantitative behaviour and will be discussed separately.
@@
The second regime arises for $\beta g>1$ and $\Delta_{0}<\tilde{\Delta}_{1}$ for some specific value $\tilde{\Delta}_{1}$, as illustrated in Fig. 3. For a given $V(\Delta)$ with $\theta=0$, there exists a continuum of self-consistent solutions of Eq. (19), corresponding to periodic orbits and limited cycles. Non-zero values of $\theta$ distort and shift the potential, introducing asymmetry. This asymmetry eliminates some previously allowed limiting cycles, since they would violate the boundedness constraint ($\Delta>\Delta_{0}$). The only remaining allowed solution is the potential minimum, which represents a non-zero fixed point of the system–a spin glass freezing state. The location of this fixed point can be tuned by adjusting $\theta$, offering a way to control the system’s behavior.
@@ VII Summary and Discussion
In our approach, we have merged the classical problem of constructing a potential well with the Waddington metaphor, extending it to a slightly deformed version when various perturbations–such as feedback mechanisms–are present in the network [8, 6, 16, 21]. These perturbations are key to analysing the stability of fixed points and the nature of bifurcations, which are critical for understanding cellular differentiation processes and disease progression, such as cancer metastasis [11, 12]. Our study begins with this motivation, focusing on the construction of a potential landscape when an epigenetic feedback term is included in the GRN. This feedback term is typically used to describe the reprogramming process in GRNs, which is crucial for cellular reprogramming or repair [8, 6, 16, 21].
@@
The complexity arises when perturbations or feedback terms are introduced into the system, adding extra dimensions to the biophysical model. This makes the construction of the potential landscape more challenging, as the additional dimensions complicate the mathematical analysis. In this work, we develop a mathematical mechanism to handle these feedback terms, which describe reprogramming in GRNs, to construct the potential landscape. This development allows us to analyse the stability of global dynamics.
We also compare the deformation of the potential landscape caused by the feedback term with the non-feedback case using the DMFT framework. The focus of our DMFT analysis is to observe how the autocorrelation function and the matching moments are affected by varying the strength of the feedback term. Our results reaveal that even small changes in the feedback term can lead to significant deformations in the potential landscape.
While the method has certain limitations–given the mathematical complexity and the length of the analysis–it provides a valuable framework for constructing a reliable version of the potential landscape of GRNs with feedback. This approach will allow us to analyse the strength of feedback and its effect on the potential surface, providing insight into various intrinsic parameters.